OpenStudy (lexi724):

Fan and Medal-- A right triangle has a hypotenuse that measures 73 mm and one side that measure 48 mm. What is the length of the remaining side of the triangle? A.55 mm B.60 mm C.65 mm D.75 mm C=73^2 B=48^2 ...right? and i have to find A

OpenStudy (bee_see):

yes

OpenStudy (lexi724):

__^2+2304=5329 5329-2304= 3025 So 55^2 answer choice A

OpenStudy (phi):

yes, and the important part is you subtract C^2 - B^2

OpenStudy (bee_see):

correct

OpenStudy (lexi724):

Thank you :)

OpenStudy (anonymous):

a^2 + 48^2 = 73^2. So do 5329 - 2304 which is 3025, and find the square root. So the answer is 55

OpenStudy (lexi724):

awesome thanks dannii